# Summarization Metrics for Spanish and Basque: Do Automatic Scores and LLM-Judges Correlate with Humans?

### v1

@@ 8 Conclusion
In this work, we present BASSE, a dataset of 2,040 automatic summaries in Basque and Spanish that we collected and annotated manually across five criteria on a 5-point Likert scale: coherence, consistency, fluency, relevance and 5W1H. The interannotator agreement study performed on a subset of triple-annotated summaries yields better scores than those found in prior large-scale summarization evaluation research. We used this dataset to perform a meta-evaluation of traditional automatic metrics and LLM-as-a-Judge models beyond English contexts. Furthermore, we compile the first automatically collected summarization dataset in Basque, which comprises 22,525 article-subhead pairs.
@@ A Annotation Guidelines
Coherence refers to the collective quality of all sentences. We align this dimension with the DUC quality question (?) of structure and coherence whereby ‘the summary should be well-structured and well-organized. The summary should not just be a heap of related information, but should build from sentence to sentence to a coherent body of information about a topic.’ To clarify, does the summary have a coherent flow of ideas and does it have connectors to explicitly define the relationships between ideas? It does not matter if there are grammatical errors or incorrect information. If the summary is simply a list of events, annotators should penalize it. Annotators do not penalize single sentences for not having connectors.
@@
B Annotation Examples
Figures 18 and 19 provide examples of low and high quality summaries in Basque and Spanish respectively, along with their corresponding average scores across all five evaluation criteria.
C Detailed Results
C.1 Interannotator Agreement
Figures 5 and 6 provide a detailed breakdown of the interannotator agreement for the final annotation round (Round #3) in Basque and Spanish respectively. We show three complementary perspectives of annotator behavior: pairwise agreement using quadratic Cohen’s $\kappa$, raw agreement percentages, and the distribution of annotations across the 5-point Likert scale.
C.2 Summarization Evaluation
C.3 Metric Correlation
Tables 5 and 6 present correlation values between automatic evaluation metrics and human annotations for Basque and Spanish respectively. For each automatic metric and human annotation criterion, both Kendall’s $\tau$ and Spearman’s $\rho$ rank correlation coefficients are reported. Rank correlation coefficients have been computed at model level.
D LLM-as-a-Judge Prompt
The following is the prompt we use for all LLM-as-a-Judge experiments, where instruction, response, reference answer, and rubric are placeholders filled with the specific information for each example.
Low-mark summary {tblr} width=colspec=*6X,row1=c,font=,row2=c,font=,rowsep=0pt,colsep=0pt,column6=bg=red9 Coh & Con Flu Rel 5W1H Avg
2.00 2.67 1.67 2.00 2.00 2.07
High-mark summary {tblr} width=colspec=*6X,row1=c,font=,row2=c,font=,rowsep=0pt,colsep=0pt,column6=bg=azure9 Coh & Con Flu Rel 5W1H Avg
4.67 4.67 4.67 4.00 5.00 4.60
Figure 3: Examples of low- and high-quality summaries of a news article in Basque. Results show the average score assigned by three annotators. Source of the article: Berria. [^18] https://www.berria.eus/euskal-herria/ehunka-herritarrek-etxebarriko-sexu-erasoa-salatu-dute_2126343_102.html
@@ D LLM-as-a-Judge Prompt
Low-mark summary {tblr} width=colspec=*6X,row1=c,font=,row2=c,font=,rowsep=0pt,colsep=0pt,column6=bg=red9 Coh & Con Flu Rel 5W1H Avg
@@
High-mark summary {tblr} width=colspec=*6X,row1=c,font=,row2=c,font=,rowsep=0pt,colsep=0pt,column6=bg=azure9 Coh & Con Flu Rel 5W1H Avg
@@
Figure 4: Examples of low- and high-quality summaries of a news article in Spanish. Results show the average score assigned by three annotators. Source of the article: El País. [^19] http://elpais.com/ccaa/2019/09/28/catalunya/1569665675_576416.html
(a) Quadratic Cohen’s $\kappa$
(b) Agreement percentage
(c) Annotation distribution across annotators and the 5-point Likert-scale
Figure 5: Interannotator agreement results for the third round (R#3) on Basque summaries.
(a) Quadratic Cohen’s $\kappa$
(b) Agreement percentage
(c) Annotation distribution across annotators and the 5-point Likert-scale
Figure 6: Interannotator agreement results for the third round (R#3) on Spanish summaries.
{tblr}
@@
[…]
Human 4.77 4.94 4.85 4.51 4.69 4.75 4.91 4.88 4.80 4.48 4.34 4.68
LLM 3.46 4.29 3.87 3.74 4.06 3.89 3.88 4.68 4.82 4.14 4.18 4.34
by model
  Claude 3.17 4.61 3.83 3.54 4.57 3.95 3.17 4.71 4.40 3.81 4.40 4.10
  GPT-4o 3.88 4.52 4.55 3.78 4.62 4.27 4.13 4.77 4.94 4.37 4.37 4.52
  Reka 3.72 4.20 4.05 4.02 4.14 4.03 4.21 4.49 4.87 4.26 4.17 4.40
  LLama 3.1 3.59 4.50 4.29 4.31 3.45 4.03 3.78 4.79 4.93 3.98 3.94 4.29
  Cmd R+ 2.95 3.65 2.64 3.07 3.53 3.17 4.10 4.66 4.94 4.29 4.02 4.40
by prompt
  Base 3.78 4.20 3.70 3.36 4.19 3.85 4.22 4.67 4.91 4.12 4.20 4.42
  Core 3.61 4.33 4.05 3.73 3.64 3.87 4.11 4.73 4.93 4.08 3.94 4.36
  5W1H 2.62 4.22 4.10 3.64 4.39 3.79 3.03 4.58 4.96 4.03 4.65 4.25
  tldr 3.82 4.42 3.64 4.24 4.01 4.03 4.14 4.75 4.47 4.35 3.94 4.33
[…]
(a) Global results, with LLM summary results averaged by model or prompt
@@
(b) Detailed results per prompting strategy across LLMs
@@
###Task Description:
###The instruction to evaluate:
{instruction}
###Response to evaluate:
{response}
###Reference Answer (Score 5):
{reference answer}
###Score Rubrics:
{rubric}
###Feedback:
Figure 7: Prompt from Prometheus-Eval used for LLM-as-a-Judge evaluation. Note that instruction, response, reference answer, and rubric are place-holders that are filled with specific information for each example.

### v2

@@ 8 Conclusion
In this work, we present BASSE, a dataset of 2,040 automatic summaries in Basque and Spanish that we collected and annotated manually across five criteria on a 5-point Likert scale: coherence, consistency, fluency, relevance and 5W1H. The interannotator agreement study performed on a subset of triple-annotated summaries yields better scores than those found in prior large-scale summarization evaluation research. We used this dataset to perform a meta-evaluation of traditional automatic metrics and LLM-as-a-Judge models beyond English contexts. Further, we share the first automatically collected summarization dataset in Basque, which comprises 22,525 article-subhead pairs.
@@
Coherence refers to the collective quality of all sentences. We align this dimension with the DUC quality question Dang (2005) of structure and coherence whereby ‘the summary should be well-structured and well-organized. The summary should not just be a heap of related information, but should build from sentence to sentence to a coherent body of information about a topic.’ To clarify, does the summary have a coherent flow of ideas and does it have connectors to explicitly define the relationships between ideas? It does not matter if there are grammatical errors or incorrect information. If the summary is simply a list of events, annotators should penalize it. Annotators do not penalize single sentences for not having connectors.
@@
[…]
Annotation Examples
Figures 22 provides an example of low- and high-quality summaries in Spanish, along with their corresponding average scores across all five evaluation criteria.
@@
Summary: Llama 3.1 70B Instruct with 5W1H prompt {tblr} width=colspec=*6X,row1=c,font=,row2=c,font=,rowsep=0pt,colsep=0pt,column6=bg=red9 Coh & Con Flu Rel 5W1H Avg
@@
Summary: GPT-4o with Base prompt {tblr} width=colspec=*6X,row1=c,font=,row2=c,font=,rowsep=0pt,colsep=0pt,column6=bg=azure9 Coh & Con Flu Rel 5W1H Avg
@@
Figure 9: Examples of low- and high-quality summaries of a news article in Spanish. Scores are averages of three annotators. Source of the article: El País. [^22] http://elpais.com/ccaa/2019/09/28/catalunya/1569665675_576416.html
@@
Interannotator Agreement
Figures 11 and 11 show interannotator agreements in terms of quadratic Cohen’s $\kappa$ for the final annotation round (Round #3) in Basque and Spanish respectively.
Figure 10: Quadratic Cohen’s $\kappa$ in Basque R#3
Figure 11: Quadratic Cohen’s $\kappa$ in Spanish R#3
@@
Summarization Evaluation
@@
[…]
Metric Correlation
Tables 6 and 7 present correlation values between automatic evaluation metrics and human annotations for Basque and Spanish respectively. For each automatic metric and human annotation criterion, both Kendall’s $\tau$ and Spearman’s $\rho$ rank correlation coefficients are reported. Rank correlation coefficients have been computed at model level.
@@
Acknowledgements.
[…]